It is believed that 55% of the eligible voters in the U.S. will vote on election day. If four eligible voters are randomly selected, what is the probability at least one will vote on election day?
Answer by stanbon(75887)

It is believed that 55% of the eligible voters in the U.S. will vote on election day. If four eligible voters are randomly selected, what is the probability at least one will vote on election day?
-------------------------
Binomial problem with n = 4 and p(vote) = 0.55 ; p(not vote) = 0.45
----------------------------
P(x>=1 votes) = 1 - P(none of 4 votes) = 1 - (0.45)^4 = 1- 0.0410 = 0.9590
============================
Cheers,
Stan H.
===============